# SplatVoxel: History-Aware Novel View Streaming without Temporal Training

### v1

@@ Abstract
We study the problem of novel view streaming from sparse-view videos, which aims to generate a continuous sequence of high-quality, temporally consistent novel views as new input frames arrive. However, existing novel view synthesis methods struggle with temporal coherence and visual fidelity, leading to flickering and inconsistency. To address these challenges, we introduce history-awareness, leveraging previous frames to reconstruct the scene and improve quality and stability. We propose a hybrid splat-voxel feed-forward scene reconstruction approach that combines Gaussian Splatting [22] to propagate information over time, with a hierarchical voxel grid for temporal fusion. Gaussian primitives are efficiently warped over time using a motion graph that extends 2D tracking models to 3D motion, while a sparse voxel transformer integrates new temporal observations in an error-aware manner. Crucially, our method does not require training on multi-view video datasets, which are currently limited in size and diversity, and can be directly applied to sparse-view video streams in a history-aware manner at inference time. Our approach achieves state-of-the-art performance in both static and streaming scene reconstruction, effectively reducing temporal artifacts and visual artifacts while running at interactive rates (15 fps with 350ms delay) on a single H100 GPU.
footnotetext: * Work done during an internship at Google.
@@ 1 Introduction
Novel view streaming with sparse views is a challenging yet essential problem with applications in augmented and virtual reality (AR/VR). The goal is to reconstruct and render temporally consistent novel views when new frames continuously come in. While recent advances in novel view synthesis, such as 3D Gaussian Splatting (3DGS) [22], enable real-time rendering, they typically require minutes of per-scene optimization using densely sampled views to construct a scene representation. This computational overhead, combined with the need for dense camera coverage, makes them impractical for online streaming reconstruction with sparse camera setups, where each frame must be processed within a reasonable time budget for video output. Feed-forward reconstruction methods [14, 5, 79] trained with large-scale scene datasets offer a more efficient alternative by enabling sparse view reconstruction and rendering without per-scene optimization. However, these methods operate on each frame independently at inference time, omitting temporal information. As a result, they usually struggle with flickering artifacts and occlusion handling.
In this paper, we tackle the problem of novel view streaming with history-awareness, leveraging information from previous frames. This is particularly crucial in sparse-view settings, where per-frame data is often insufficient and occlusions are common. By incorporating history, we can enhance reconstruction quality and mitigate temporal flickering artifacts, significantly improving the user experience. However, working on multi-view dynamic scene reconstruction presents its own challenges. Large-scale, multi-view video datasets suitable for training are very limited, and efficiently modeling temporal correspondences is nontrivial. To this end, we present a novel feed-forward scene reconstruction method that effectively leverages historical information for streaming reconstruction without requiring training on the currently limited multi-view video datasets. Our approach introduces a hybrid Splat-Voxel representation that combines the advances of Gaussian Splatting [22] and an efficient sparse voxel transformer. Specifically, we adopt Gaussian Splatting as a moveable point-based representation, enabling real-time splatting-based rendering and facilitating deformation across frames. We warp the historical feed-forward reconstruction of Gaussian Splatting using an efficient motion uplifting algorithm, which leverages existing 2D motion priors from pretrained models [10, 64]. To integrate historical information carried from warped Gaussian primitives with new observations from the current frame, we introduce a sparse voxel transformer with a coarse-to-fine voxel representation, enabling effective information fusion within a temporally shared Euclidean space. Our method achieves state-of-the-art performance in terms of rendering quality and temporal consistency on several multi-view video datasets under the sparse camera setup, and can reach interactive rate (15fps with 350ms delay) under a single H100 GPU. In summary, the main contributions of our paper are as follows:
  • We propose a hybrid Splat-Voxel feed-forward reconstruction framework that leverages historical information to enable novel view streaming, without relying on multi-view video datasets for training.
  • We develop an efficient sparse voxel transformer with a coarse-to-fine voxel representation, outperforming existing feed-forward Gaussian splatting methods.
  • Experiment results demonstrate that our proposed framework enhances novel view synthesis for streaming scene reconstruction with better visual quality and reduced temporal artifacts through history-aware modeling.
@@ 2 Related Work
Image-based rendering approaches for novel view synthesis (NVS) predict the appearance of a scene from novel viewpoints given a set of input images. A key design decision is the choice of 3D scene representation to establish correspondences and capture structure. Popular representations include multi-plane images [13, 82, 4, 71, 65, 72, 27], neural fields [40, 56], voxel grids [61, 41, 50], and Gaussian splatting [22]. NVS methods can be broadly divided into those that optimize scene representation at test time and those that directly predict it through a feed-forward network [67, 59, 52, 78, 5, 12]. Recently, transformers have been used to reason about scenes without explicit 3D bias [20, 54, 53] or in combination with differentiable renderers in a hybrid approach [79, 17]. Our method follows the latter paradigm, leveraging a transformer backbone that integrates voxel grids and 3D Gaussians for scene representation while operating in a data-driven feed-forward manner to achieve streaming applications.
@@
Handling dynamic motions in novel view synthesis (NVS) is challenging due to occlusions and the need for temporal consistency. Dynamic scene reconstruction is typically addressed in multi-view setups with fixed cameras [75, 57, 73, 8, 38] or monocular settings where both the camera and scene move [43, 28, 74, 29, 68]. Most methods optimize on the entire video sequences, relying on learned or hand-crafted priors such as optical flow [64, 36], tracking [10], and depth estimation [44, 76, 48, 49] for supervision [28, 74]. To regularize motion, these methods often incorporate basis functions [29, 68, 30] or motion priors [47], while others explicitly model static and dynamic components separately [21, 32, 35]. These techniques support applications such as animation [63, 18, 47], tracking [68, 39], and semantic understanding [33]. Some methods accelerate reconstruction by preserving past reconstructions for faster optimization [69, 25, 62], but they remain computationally expensive and impractical for real-world, sparse-view streaming scenarios. Our method uses pretrained models (specifically, point tracking) for motion priors. However, unlike typical methods that optimize on the entire video sequence, our model operates causally and online, using only previously received frames.
Online 3D Reconstruction
Feed-forward approaches achieve online 3D reconstruction by predicting novel views in a single pass without per-scene optimization. However, many rely solely on the current frame [14, 34, 51, 31], neglecting temporal information, which leads to flickering and occlusion artifacts. Some methods [51, 31] train on large-scale dynamic video datasets, often incorporating synthetic data due to the scarcity of real-world multi-view videos. However, they assume access to full sequences at inference, limiting their applicability in real-time settings. Our approach bridges these gaps by combining the efficiency of feed-forward inference with history-awareness. Unlike optimization-based methods, we do not require future frames, and unlike standard feed-forward models, we leverage past frames to improve temporal consistency. By generalizing from static scene training to dynamic video inference without additional temporal supervision, our method enables efficient, real-time novel view streaming from sparse multi-view inputs.
Table 1: Method Comparison. Unlike existing baselines, our framework uniquely supports the three key requirements of the novel view streaming task targeted in this paper: fast reconstruction with interactive rate , sparse view inputs, and history awareness. In particular, our streaming approach only uses past frames, in constrast to DeformableGS [77] and DeformableNerf [43] that optimize over $\textit{all}^{*}$ frames. The history-awareness in our model mitigates temporal flickering and occlusion artifacts, and is also able to support unlimited sequence lengths.
Method  Interactive Speed  Sparse-view Input  History Awareness
NeRF [40] $\times$ $\times$ $\times$
InstantNGP [41] $\times$ $\times$ $\times$
3DGS [22] $\times$ $\times$ $\times$
[…]
DeformableNerf [43] $\times$ $\checkmark^{*}$
[…]
GS-LRM [79] $\times$
Quark [14] $\times$
Ours
Figure 2: Method Overview. Our hybrid SplatVoxel model first extracts input image features using a multi-view transformer, which outputs pixel-aligned Gaussian splats for each input image with associated features. The splat features are then deposited onto a coarse-to-fine voxel grid using the decoded position, and a secondary sparse voxel transformer processes the grid features to output final Gaussian parameters. To merge history, we compute the triangulated scene flow from the input views and perform keypoint guided deformations. These deformed splats can be treated identically to the input-aligned splats, and be similarly deposited into voxel grid to merge the previous state with the current state. Our system is trained only on static scenes, and can generalize to dynamic scenes during inference.
3 Online Novel View Streaming
Given sparse (e.g., 2–4 views) RGB video streams from fixed and calibrated cameras, we aim to achieve online novel view synthesis, using only temporal information from past frames. The key objectives for our approach include efficient reconstruction (i.e. $>10$ fps interactive rate) and effective reuse of past information. History awareness is essential for our task given the sparse coverage of the input cameras, which can lead to artifacts such as temporal flickering and degraded reconstruction in occluded regions. Rather than processing each frame independently, using information from prior frames improves the per-frame reconstruction quality and ensures temporal consistency.
@@ 3 Online Novel View Streaming
Optimization-based methods [40, 2, 41, 22, 73] are not well-suited for this task, as they require extensive optimization over entire video sequences. Additionally, they perform poorly with sparse views and face challenges in representing long video sequences due to memory and capacity constraints. Recent feed-forward reconstruction methods [5, 6, 79, 14] have shown promise in efficient sparse-view reconstruction of static scene by utilizing large networks trained on large-scale multi-view datasets. However, their lack of explicit temporal modeling makes them prone to temporal flickering and occlusions in video inputs, resulting in suboptimal performance for our novel view streaming task.
Table 1 highlights the differences among these prior methods, none of which address novel view streaming or incorporate historical information. Integrating history-awareness into existing frameworks is particularly challenging. A major obstacle in training history-aware feed-forward networks is the limited availability of high-quality multi-view video datasets. Moreover, naive approaches, such as continuously warping historical scene reconstructions to later frames, fail to capture newly appearing objects and changes in appearance—not to mention that accurate 3D scene motion estimation from sparse multi-view RGB inputs remains challenging and prone to accumulated errors, ultimately degrading reconstruction fidelity.
Our Solution
To address these challenges, we propose a novel feed-forward reconstruction framework that generalizes to temporal video streams at inference time, despite being trained solely on multi-view static scenes. Our approach leverages a hybrid point-voxel representation, combining the flexibility of point-based methods with the structured spatial organization of voxels, drawing inspiration from hybrid Lagrangian-Eulerian methods [19] commonly used in physics simulation.
Our two-stage pipeline first constructs a movable point-based (Lagrangian) representation, which is then transferred into a voxel-based (Eulerian) representation. This design naturally extends static scene reconstruction to history-aware dynamic scene reconstruction, where historical information is retained within the Lagrangian component and continuously integrated with new data on the global Eulerian grid, following the principles of hybrid Lagrangian-Eulerian methods.
For our novel view streaming task, we employ Gaussian Splatting [22] as the Lagrangian component and a coarse-to-fine sparse voxel grid as the Eulerian component. A multi-view transformer learns to generate Lagrangian Gaussian primitives from multi-view RGB images, while a voxel transformer fuses information within the Eulerian grid in latent space. After training on a large-scale static scene dataset, our method enables online novel view streaming without additional temporal training. This is achieved by warping Gaussian primitives using an embedded deformation graph [60] derived from uplifted 2D tracks [10] and integrating the warped primitives with the current frame using the voxel transformer.
4 Method
In this section, we describe our SplatVoxel feed-forward scene reconstruction framework (Sec 4.1), and how it can generalize at inference time for zero-shot history-aware 4D novel view streaming (Sec 4.2).
4.1 Efficient SplatVoxel Transformer
Our two-stage feed-forward reconstruction model fuses information across 2D and 3D domains by first using a multi-view transformer to process 2D images into Gaussian splat primitives (we refer to a 3D Gaussian primitive as a “splat” for brevity here and in the following sections). We then deposit the splats on a voxel grid and train a sparse coarse-to-fine voxel transformer to refine the Gaussian primitives.
4.1.1 Multi-view Transformer
Given $N$ RGB images of resolution $H\times W$ with known camera parameters, we extract multi-view latent features and train a transformer following GS-LRM [79]. We concatenate the RGB value with per-pixel Plücker ray embeddings [45]. Next, we extract features from non-overlapping $p\times p$ image patches [11] using a single-layer CNN, resulting in patch features with channel dimension of $C$. The resulting multi-view image patch features are concatenated into a set of $N\times HW/p^{2}$ tokens, which we denote as $\{\mathbf{X}\}$. The multi-view (MV) transformer [66] then processes image tokens $\mathbf{X}\}$ to capture the underlying geometric and appearance, producing a set of latent features $\{\mathbf{Z}\}$:
$\centering\{\mathbf{Z}\}=\mathrm{Transformer}_{MV}(\{\mathbf{X}\})\@add@centering$ (1)
We unpatchify the processed latent features to Gaussian primitives with a linear layer applied to each latent:
$\centering\{\mathbf{G}_{k},\mathbf{F}_{k}\}=\mathrm{Linear}(\{\mathbf{Z}\})\@add@centering$ (2)
where $\mathbf{G}_{k}$ is a Gaussian primitive aligned with each input pixel ($k=1...,NHW$) which can be rendered using Gaussian splatting [22], and $\mathbf{F}_{k}$ is an associated feature vector for each Gaussian primitive.
4.1.2 Splat-to-Voxel Transfer
While previous feed-forward Gaussian Splatting methods [6, 79, 5] typically represent the scene by simply concatenating per-pixel splats predicted from different views, we propose a 3D-aware refinement step on splats $\mathbf{G}_{k}$ and associated features $\mathbf{F}_{k}$ by leveraging an intermediate voxel-based representation. This Splat-to-Voxel Transfer is achieved by first performing Splat Depositing onto a voxel grid, followed by Splat Fusion in Euclidean space to produce the input for our sparse voxel transformer.

### v2

# Learning Efficient Fuse-and-Refine for Feed-Forward 3D Gaussian Splatting

@@ Abstract
Recent advances in feed-forward 3D Gaussian Splatting have led to rapid improvements in efficient scene reconstruction from sparse views. However, most existing approaches construct Gaussian primitives directly aligned with the pixels in one or more of the input images. This leads to redundancies in the representation when input views overlap and constrains the position of the primitives to lie along the input rays without full flexibility in 3D space. Moreover, these pixel-aligned approaches do not naturally generalize to dynamic scenes, where effectively leveraging temporal information requires resolving both redundant and newly appearing content across frames. To address these limitations, we introduce a novel Fuse-and-Refine module that enhances existing feed-forward models by merging and refining the primitives in a canonical 3D space. At the core of our method is an efficient hybrid Splat-Voxel representation – from an initial set of pixel-aligned Gaussian primitives, we aggregate local features into a coarse-to-fine voxel hierarchy, and then use a sparse voxel transformer to process these voxel features and generate refined Gaussian primitives. By fusing and refining an arbitrary number of inputs into a consistent set of primitives, our representation effectively reduces redundancy and naturally adapts to temporal frames, enabling history-aware online reconstruction of dynamic scenes. Trained on large-scale static scene datasets, our model learns an effective global strategy to process around 200k primitives within 15ms and significantly enhances reconstruction quality compared to pixel-aligned reconstruction approaches. Without additional training, our model generalizes to video by fusing primitives across time, yielding a more temporally coherent result compared to baseline methods with graceful handling of occluded content. Our approach achieves state-of-the-art performance in both static and streaming scene reconstructions while running at interactive rates (15 fps with 350ms delay) on a single H100 GPU.
@@ 1 Introduction
Novel view synthesis from sparse views is a core problem for augmented and virtual reality (AR/VR) applications, where practical constraints require the reconstruction to be fast and accurate for a positive user experience. Recent feed-forward 3D Gaussian Splatting methods xu2024grm; charatan2024pixelsplat; zhang2025gslrm trained with large-scale scene datasets offer efficient solutions for the sparse-view reconstruction problem. The key idea of these approaches is to learn a large feed-forward network, typically Transformer-based, to predict Gaussian primitives and use splatting as a fast renderer kerbl20233dgs.
The common approach taken by these methods is to regress per-pixel Gaussian primitives that lie along the input camera rays. However, this leads to redundancy in the representation when input views have overlapping content, and can also lead to incomplete geometry with occluded content not visible from the input views. This per-pixel representation also does not naturally generalize to dynamic scenes, where redundant primitives between frames should be merged and new primitives should emerge to capture newly appearing content.
To address these limitations, we introduce a novel Fuse-and-Refine module that processes and merges Gaussian primitives in a canonical 3D space. Our approach enables more flexible primitive placement and better density control, leading to improved reconstruction quality in static scenes. It also greatly facilitates streaming reconstruction in dynamic scenes by avoiding error accumulation and redundant primitives over time, while better handling temporal occlusions and disocclusions through the use of historical information. We note that past approaches to merging primitives have relied on heuristics hierarchicalgaussians24 that typically degrade reconstruction quality and require additional time-consuming gradient descent optimization steps. In contrast, our learning-based solution leverages a feed-forward Transformer trained on large-scale scene datasets combined with the structured nature of voxels as an intermediate scaffold to efficiently fuse and refine Gaussian primitives.
We design a hybrid Splat-Voxel model that proceeds as follows: given a set of initial Gaussian primitives from existing pixel-aligned feed-forward models, we first perform a Splat-to-Voxel Transfer to aggregate local features from Gaussian primitives into a voxel representation, and then apply a Transformer to generate refined Gaussian primitives from voxel features. Our voxel representation is specifically designed to ensure both efficiency and high-quality output. We use a coarse-to-fine voxel hierarchy to balance detail-preservation and computational overhead, operating a sparse voxel transformer on features at the coarse resolution while refining the primitives at a high resolution. Our Fuse-and-Refine module can merge 200K Gaussian primitives from existing feed-forward 3D Gaussian models in just 15 ms, while delivering a PSNR improvement of around 2 dB.
We train our model exclusively on static datasets, but our model also enables history-aware streaming reconstruction of sparse-view dynamic scenes without additional training or modifications. This is advantageous as multiview video datasets remain limited in both size and diversity, prohibiting large-scale training. For dynamic scenes, we first warp primitives from a past frame to the current frame, filtering out those with large warping errors, and deposit primitives from both the past and current frame onto our Splat-Voxel representation. The resulting refined primitives significantly improve the temporal consistency of existing feed-forward models, while avoiding redundant primitives and mitigating error accumulation. Our method achieves state-of-the-art performance on several multi-view video datasets under the sparse camera setup, and can reach interactive rates (15fps with 350ms delay) under a single H100 GPU.
In summary, the main contributions of our paper are as follows:
  • We introduce a novel Fuse-and-Refine module that enhances existing feed-forward 3D Gaussian Splatting methods and extends them to history-aware streaming reconstruction in a zero-shot manner.
  • We propose a hybrid Splat-Voxel representation that combines a coarse-to-fine voxel hierarchy with a sparse voxel Transformer, enabling effective learning of the Fuse-and-Refine module on large-scale static scene datasets.
  • Extensive experimental results demonstrate that our approach achieves state-of-the-art performance on both static and streaming-based dynamic scene reconstruction, while maintaining interactive runtimes.
@@ 2 Related Work
A wide range of methods have been developed over the past decades for novel view synthesis (NVS). A key design decision is the choice of 3D scene representation to establish correspondences and capture structure. Popular representations include multi-plane images flynn2016deepstereo; flynn2019deepview; broxton2020immersive; wiles2020synsin; tucker2020single; wizadwongsa2021nex; li2020crowdsampling, neural fields mildenhall2021nerf; sitzmann2019scene, voxel grids sun2022direct; muller2022instantngp; reiser2024binary, and point clouds xu2022pointnerf; kerbl20233dgs. In this work, we adopt 3D Gaussian Splatting kerbl20233dgs as our final rendering representation due to its ability to achieve high-quality and real-time consistent novel view synthesis.
Existing NVS methods can be broadly divided into those that optimize scene representation at test time and those that directly predict it through a feed-forward network wang2021ibrnet; suhail2022gpnr; riegler2020free; yu2021pixelnerf; du2023learning; jin2024lvsm; sajjadi2022scene; rombach2021geometry; hong2023lrm. The combination of Transformers and 3D Gaussian Splatting has emerged as an efficient approach to predict pixel-aligned Gaussian splatting in a data-driven, feed-forward manner zhang2025gslrm; chen2024mvsplat; charatan2024pixelsplat; zhang2025transplat; tang2024hisplat; wei2025omni. However, pixel-aligned prediction design often results in suboptimal scene reconstruction due to limited flexibility in accurately positioning primitives in 3D space. To address this limitation, several works have leveraged feed-forward networks operating on 3D representations, such as graph zhang2024gaussianGNN, voxel miao2025evolsplat; ren2024scube; LaRa, or point cloud chen2024splatformer, to process 3D Gaussian Splatting. Yet, extending these representations to dynamic scenes remains challenging and largely unexplored. To this end, we propose a hybrid point-voxel representation that can generalize to dynamic scenes during inference while trained only on static scenes. Our design shares a similar spirit with voxel-based point cloud processing networks zhou2018voxelnet; liu2019pointvoxelCNN; mao2021voxeltransformer, which leverage voxel structures for efficient point cloud processing. However, existing methods li2024bevformer typically rely on autonomous-driving–specific projections (e.g., bird’s-eye view) to handle large-scale scenes. In contrast, we construct a cost volume suitable for general unbounded scenes and introduce a coarse-to-fine voxel hierarchy that enables global attention at the coarsest level.
@@
Handling dynamic motions in novel view synthesis (NVS) is challenging due to occlusions and the need for temporal consistency. Dynamic scene reconstruction is typically addressed in multi-view setups with fixed cameras xu2024longvolcap; stearns2024gsmarble; wu20244d; Collet2015streamingvideo; lombardi2019neuralvolume; SplatFields or monocular settings where both the camera and scene move park2021nerfies; li2021neural; xian2021space; li2023dynibar; wang2024shape. Most methods optimize on the entire video sequence, relying on learned or hand-crafted priors such as optical flow teed2020raft; lin2022occlusionfusion, tracking doersch2023tapir, and depth estimation piccinelli2024unidepth; yang2024depth; ranftl2020towards; ranftl2021vision for supervision li2021neural; xian2021space. Our method leverages pretrained models for point tracking and optical flow as motion priors and operates causally and online, using only previously received frames without requiring access to the full video.
Streaming reconstruction Collet2015streamingvideo presents a significant challenge in dynamic scene reconstruction, aiming to enable online novel view synthesis from potentially unbounded video streams for real-world applications. Several methods wangneus2; li2022streamingNeRF; sun20243dgstream; gao2024hicom have explored the use of historical information to accelerate optimization, but remain computationally expensive and depend on dense multi-view inputs. Feed-forward models are well-suited for sparse multi-view streaming scene reconstruction. However, many of these methods can only take a single frame as input flynn2024quark; lin2022efficient; ren2025l4gm; zhang2025gslrm; charatan2024pixelsplat, ignoring temporal information and consequently suffering from flickering and occlusion artifacts. Moreover, the scarcity of real-world multi-view dynamic datasets hinders the ability to train feed-forward dynamic models ren2025l4gm; liang2024feed for real-world streaming applications. By generalizing from static scene training to dynamic video inference without requiring additional temporal data, our method enables efficient, history-aware novel view synthesis for streaming scenarios.
3 Preliminaries: Feed-forward 3D Gaussian Splatting
3D Gaussian Splatting (3DGS) kerbl20233dgs
3DGS represents a scene as a set of anisotropic 3D Gaussian primitives, each defined by its position $\boldsymbol{\mu}_{i}\in\mathbb{R}^{3}$, opacity $\alpha_{i}\in\mathbb{R}$, quaternion rotation $\mathbf{r}_{i}\in\mathbb{R}^{4}$, scale $\mathbf{s}_{i}\in\mathbb{R}^{3}$, and spherical harmonic coefficients $\mathbf{c}_{i}\in\mathbb{R}^{S}$, where $S$ denotes the number of coefficients used for modeling view-dependent color. The rendered color at pixel $\mathbf{p}$, denoted as $C(\mathbf{p})$, is computed via splatting-based rasterization, where each 3D Gaussian is projected to 2D screen space and blended using front-to-back alpha compositing:
$C(\mathbf{p})=\sum_{i=1}^{K}\mathbf{c}_{i}\alpha_{i}\mathcal{G}_{i}^{\text{2D}}(\mathbf{p})\prod_{j=1}^{i-1}\left(1-\alpha_{j}\mathcal{G}_{j}^{\text{2D}}(\mathbf{p})\right)$ (1)
Here, $\mathcal{G}_{i}^{\text{2D}}(\mathbf{p})$ denotes the 2D projection of the $i$-th 3D Gaussian at pixel $\mathbf{p}$.
Feed-forward Reconstruction Model
To predict the 3D Gaussian primitives in a feed-forward manner from sparse-view images, we adopt the multi-view Transformer architecture from GS-LRM zhang2025gslrm. Given $N$ RGB images of resolution $H\times W$ with known camera intrinsics and extrinsics, we extract per-pixel features by concatenating RGB values with Plücker ray embeddings plucker1865xvii. These features are grouped into non-overlapping $p\times p$ image patches dosovitskiy2020vit and encoded using a shallow CNN to obtain patch features of channel dimension $C$. The patch features from all $N$ views are then flattened into a sequence of $N\times\frac{HW}{p^{2}}$ tokens, denoted as $\{\mathbf{X}\}$. These tokens are processed by a multi-view Transformer vaswani2017transformer to jointly model geometric structure and appearance across views, producing latent features $\{\mathbf{Z}\}$:
$\{\mathbf{Z}\}=\mathrm{Transformer}_{MV}(\{\mathbf{X}\})$ (2)
Finally, a linear projection is applied to each latent token to produce an initial set of Gaussian primitives and associated feature vectors:
$\{\mathbf{G}_{k},\mathbf{F}_{k}\}=\mathrm{Linear}(\{\mathbf{Z}\})$ (3)
Here, $\mathbf{G}_{k}$ represents the parameters of a 3D Gaussian primitive aligned with the $k$-th image patch, which can be directly rendered using Gaussian Splatting kerbl20233dgs, while $\mathbf{F}_{k}$ denotes an additional feature vector used for the subsequent learning-based fusion and refinement.
Figure 1: Method Overview. Our Fuse-and-Refine module takes as input Gaussian primitives produced by existing feed-forward models, either from the current frame or warped from previous reconstructions, and produces fused and refined primitives that improve scene reconstruction. The input primitives are first deposited into a high-resolution voxel grid with a splat-to-voxel Transfer strategy, which is then adaptively sparsified to construct a coarse-to-fine voxel hierarchy. A sparse voxel transformer is applied at the coarse level to capture global context, and new primitives are subsequently generated at the high-resolution level.
4 Learning to Fuse and Refine
Given an arbitrary set of initial Gaussian primitives, for instance those from multi-view images (as in Sec. 3) or from previous time steps (as we will describe in Sec. 5), we aim to fuse nearby primitives and produce refined primitives that improve novel view synthesis results.
@@ 4 Learning to Fuse and Refine
Preserving the fidelity of millions of Gaussian primitives while merging them is highly non-trivial, with several factors contributing to the challenge. First, maintaining the original 3D Gaussian distributions without divergence is fundamentally infeasible when combining multiple kernels. This issue is further exacerbated in the context of sparse-view novel view synthesis, where the fusion problem is inherently ill-posed. In addition to merging the Gaussian distribution determined by primitive mean and covariance, other rendering attributes like color, opacity, scale, and rotation must also be reconciled with limited supervision from sparse 2D observations. As a result, heuristic approaches hierarchicalgaussians24 often lead to degraded reconstruction quality.
Solution
Instead, we propose an efficient learning-based approach that not only merges 3D primitives, but also further refines them through a feed-forward Transformer. Trained on large-scale scene data, our model learns a global strategy to fuse input primitives into a refined set of splats in just 15 ms, significantly improving reconstruction performance. The proposed learning framework adopts a hybrid Splat-Voxel representation, where an intermediate voxel grid serves as a structured spatial proxy for aggregating and distributing splats. A voxel Transformer operates in latent space to fuse information in a canonical 3D space and produce refined splats. An overview of the proposed Fuse-and-Refine module is illustrated in Fig. 1.
4.1 Splat-to-Voxel Transfer
Starting with the predicted Gaussian primitives $\mathbf{G}_{k}$ and their associated image features $\mathbf{F}_{k}$ from the multi-view Transformer (Eqn. 2), we first operate a Splat-to-Voxel Transfer procedure to convert the point-based representation to a structured voxel representation. Implementation details, including the choice of kernel function $\mathcal{K}$ and other parameters, are provided in the supplementary material (Sec. A).